# Integrated Sensing and Communication for Low-Altitude Security

### v1

@@ I Introduction
With the completion of the first standard (3GPP Release 18) of 5G-Advanced in 2024, the global evolution of wireless networks toward 6G has formally entered a new phase. Within this technological transition, integrated sensing and communication (ISAC) has emerged as a key enabling technology for both 5G-Advanced and 6G. By enabling the reuse of network infrastructure and spectrum resources, ISAC achieves deep synergy and significant performance enhancement between communication and radar sensing functions [1, 2], garnering extensive attention from both academia and industry.
@@
Against this backdrop, ISAC technology offers a highly promising innovative approach to addressing the challenges of low-altitude security (LAS) governance. Essentially, ISAC aims to upgrade cellular networks from mere communication pipelines to intelligent infrastructure with integrated environmental sensing capabilities [7]. This provides a novel paradigm for tackling the LAS predicament of being undetected, misidentified and uncontrollable: leveraging ubiquitous mobile communication base stations (BSs) as distributed sensing nodes, it is expected to achieve continuous, high-resolution and intelligent situational awareness of low-altitude airspace at an affordable cost, and integrate with communication capabilities to establish a closed-loop security management system covering sensing, cognition and decision-making [4].
This article systematically reviews the ISAC technical system for LAS. It first analyzes the unique security challenges in low-altitude scenarios. Subsequently, it prospects the empowerment of ISAC for LAS against each challenge. Finally, it discusses the current open problems and challenges. To the best of the authors’ knowledge, this is the first review article focusing on ISAC for LAS.
@@ II-B Key Security Challenges
More importantly, systems cannot distinguish whether a flight activity is an approved legal operation, unintentional unreported flight (grey flight), or malicious unauthorized intrusion (black flight). They lack the capability of real-time analysis based on multi-dimensional data for high-risk behavioral patterns (e.g., abnormal loitering and formation gathering of targets around sensitive areas), failing to achieve the cognitive leap from trajectory tracking to intention understanding and threat determination.
@@ III ISAC-Enabled LAS
Addressing the security challenges enumerated in Section II-B, this section elaborates on the approaches and advantages of ISAC in empowering LAS.
@@ III-B Intelligent Feature Extraction and Intention Understanding
For high-level intention understanding, it is necessary to construct a behavior analysis engine based on temporal deep learning models or knowledge graphs to model targets’ continuous trajectories, speed changes, and regional dwell patterns. Integrating context provided by geographic information systems (e.g., distribution of sensitive facilities, airspace structure), the system can identify abnormal behavioral patterns such as loitering, approaching reconnaissance, and formation gathering, and perform probabilistic prediction of potential intentions (e.g., straying, snooping, malicious intrusion) and threat level assessment.
@@ III-C Real-Time Collaborative Situation Sharing and Decision Support
To address the conflict resolution challenge, as shown in Fig. 4, the constructed ISAC network can generate a full-domain real-time situational map, which is converted into shareable public information products via low-latency communication, thereby promoting the establishment of a new security paradigm driven by network-wide perception and aerial distributed collaboration [5].

### v2

@@ I Introduction
With the completion of the first standard (3GPP Release 18) of 5G-Advanced in 2024, the global evolution of wireless networks toward 6G has formally entered a new phase. Within this technological transition, integrated sensing and communication (ISAC) has emerged as a key enabling technology for both 5G-Advanced and 6G. By enabling the reuse of network infrastructure and spectrum resources, ISAC achieves deep synergy and significant performance enhancement between communication and radar sensing functions [22], garnering extensive attention from both academia and industry.
@@
Against this backdrop, ISAC technology offers a highly promising innovative approach to addressing the challenges of low-altitude security (LAS) governance. Essentially, ISAC aims to upgrade cellular networks from mere communication pipelines to intelligent infrastructure with integrated environmental sensing capabilities [14, 7]. This provides a novel paradigm for tackling the LAS predicament of being undetected, misidentified and uncontrollable: leveraging ubiquitous mobile communication base stations (BSs) as distributed sensing nodes, it is expected to achieve continuous, high-resolution and intelligent situational awareness of low-altitude airspace at an affordable cost, and integrate with communication capabilities to establish a closed-loop security management system covering sensing, cognition and decision-making [4].
ISAC has emerged as a key technical approach to bolstering the development of the LAE. Current research has been conducted from multiple dimensions. In terms of waveform and signal processing, research focuses on designing waveforms with both high communication reliability and high sensing resolution. For instance, the work of [4] proposed an adaptive waveform, which optimized the bit error rate and sensing sidelobe level by integrating coding and index modulation. The work of [16] improved the accuracy and robustness of UAV parameter estimation via tensor decomposition and multi-station data fusion.
In network architecture and resource optimization, research focuses on improving the overall system efficiency via collaborative and intelligent technologies. Ref. [10] introduced a hybrid reconfigurable intelligent surface to compensate for the elevation angle limitation of BS antennas and improved the sensing signal-to-noise ratio for high-altitude targets. Ref. [1] balanced sensing and communication performance through the joint design of collaborative beamforming and UAV trajectories.
Several studies propose novel system frameworks and performance metrics. Ref. [3] presentd a full-duplex ISAC system prototype, enabling simultaneous co-frequency sensing and communication via interference cancellation. Ref. [5] modeled low-altitude surveillance as a compressed sensing imaging problem and introduces physical embedding learning to mitigate off-grid errors. Ref. [17] innovatively proposed the metric of sensing capacity to evaluate the upper limit of BSs’ ability to detect multiple targets simultaneously.
These efforts have advanced the technological development of ISAC in low-altitude environments and preliminarily verified its potential to improve sensing accuracy, communication efficiency and resource utilization. However, most existing studies focus on enhancing the communication-sensing performance of ISAC systems themselves or address specific technical challenges (e.g., interference management, trajectory design). Although several studies mention the importance of LAE [6, 2], their analyses are often embedded in the framework of ISAC performance optimization, lacking a comprehensive sorting and in-depth analysis of security challenges, ISAC enabling paths and existing problems from the systematic perspective of LAE governance.
This article thus focuses on the ISAC technical system for low-altitude security. First, Section II analyzes the unique security challenges of low-altitude scenarios. Section III then prospects ISAC-enabled LAE for each challenge. Section IV discusses the current open problems and challenges. Section V concludes the article. To the authors’ knowledge, this is the first review article focusing on ISAC for LAS.
@@ II-B Key Security Challenges
More importantly, systems cannot distinguish whether a flight activity is an approved legal operation, unintentional unreported flight (grey flight), or malicious unauthorized intrusion (black flight). They lack the capability of real-time analysis based on multi-dimensional data for high-risk behavioral patterns (e.g., abnormal loitering and formation gathering of targets around sensitive areas [21]), failing to achieve the cognitive leap from trajectory tracking to intention understanding and threat determination.
@@ III ISAC-Enabled LAS
Addressing the security challenges enumerated in Section II, this section elaborates on the approaches and advantages of ISAC in empowering LAS.
@@ III-B Intelligent Feature Extraction and Intention Understanding
For high-level intention understanding, it is necessary to construct a behavior analysis engine based on temporal deep learning models or knowledge graphs to model targets’ continuous trajectories, speed changes, and regional dwell patterns [12]. Integrating context provided by geographic information systems (e.g., distribution of sensitive facilities, airspace structure), the system can identify abnormal behavioral patterns such as loitering, approaching reconnaissance, and formation gathering, and perform probabilistic prediction of potential intentions (e.g., straying, snooping, malicious intrusion) and threat level assessment.
@@ III-C Real-Time Collaborative Situation Sharing and Decision Support
To address the conflict resolution challenge, as shown in Fig. 4, the constructed ISAC network can generate a full-domain real-time situational map, which is converted into shareable public information products via low-latency communication, thereby promoting the establishment of a new security paradigm driven by network-wide perception and aerial distributed collaboration [16, 11].